OpenStudy (anonymous):

m^2 over m^15 divide and simplify

OpenStudy (ksaimouli):

m^2-15 m^-13

OpenStudy (saifoo.khan):

\[\frac{m^2}{m^{15}}=\frac{1}{m^{15-2}}=\frac{1}{m^{13}}\]

OpenStudy (ksaimouli):

1/m^13